3 If sin(0) = and 0 is in the 3rd quadrant, find the exact value for cos(0). cos(0)
State the quadrant in which 6 lies. tan(0) > 0, csc(0) < 0 OTI OMI OIV
Answer #1

Given sino = 910 O = sint (233 0 - 180 +36.87 = 216.87 Now Go= G) (216 87 ) - -0.8 Z a 2 Gimn tanoro CSC10) <0. in 3rd quadra